OpenStudy (anonymous):

Find cos 36 with Trigonometry

OpenStudy (anonymous):

or you can find cos 36 with a calculator :D

OpenStudy (anonymous):

you have to do it from scratch or can we use the value of \(\sin 18^\circ \)?

OpenStudy (anonymous):

do not use known values

OpenStudy (anonymous):

http://mathforum.org/library/drmath/view/52139.html

OpenStudy (callisto):

let x = 36 5x = 180 3x = 180 - 2x sin 3x = sin2x 3 sinx - 4sin^3 x = 2sinx cox sinx =/= 0 3- 4 sin^2 x = 2cosx 3- 4 ( 1-cos^2 x) = 2 cos x 4 cos^2 x - 2cosx -1 =0 \[cosx = \frac{-(-2) \pm \sqrt{(-2)^2-4(1)(-1)}}{2(4)}\] \[cos36 = \frac{2 \pm \sqrt{20}}{8} = \frac{1 \pm \sqrt{5}}{4} \]

OpenStudy (callisto):

assume that 36 is in degrees...

OpenStudy (anonymous):

what is "sinx =/= 0" ?

OpenStudy (anonymous):

How did 3x=180-2x suddenly turn into sin 3x = sin 2x? I'm sorry if I'm a bit slow :D

OpenStudy (anonymous):

Taking sine of the both sides.

OpenStudy (callisto):

@im2bad sinx≠0 since sinx = sin 36 ≠0

OpenStudy (anonymous):

sin(pi-a)=sin a

OpenStudy (callisto):

sin (180 -2x) = sin 2x

OpenStudy (aravindg):

nic work callisto

OpenStudy (anonymous):

Right. Thank you.

OpenStudy (callisto):

sin 3x = sin (2x+x) = sin2x cosx + cos2x sinx = (2sinx cosx) cosx + (1-2sin^2x)sinx = 2sinx cos^2x + (1-2sin^2x)sinx = 2sinx (1-sin^2 x) + (1-2sin^2x)sinx = 2sinx - 2sin^3 x + sinx - 2sin^3 x = 3sinx - 4sin^3 x

OpenStudy (callisto):

Ahh.. wait :| One more step to go since cos 36>0 \[cos36 =\frac{1+\sqrt{5}}{4}\]

OpenStudy (anonymous):

how about this \[\cos 36=x \ \ then \ \ \cos 72=2x^2-1 \\ \cos 36-\cos 72=2 \ \sin 54 \ \sin 18=2 \cos 36 \ \cos 72 \\ x-2x^2+1=2x(2x^2-1) \\ 4x^3+2x^2-3x-1=0 \\ (4x^2-2x-1)(x+1)=0 \\ \cos 36 \neq -1 \\ 4x^2-2x-1=0 \\ x=\frac{1+\sqrt{5}}{2}\]

OpenStudy (anonymous):

err \[x=\frac{1+\sqrt{5}}{4}\]

OpenStudy (callisto):

I learnt another way to solve it! Amazing!!!! ><

OpenStudy (anonymous):

i forgot to medal u :)